OpenStudy (anonymous):

Fool's problem of the day! A Fascinating counting problem, In how many ways can we paint a staircase of \( 50 \) steps with two colors green and red such that green steps should never succeed each other? PS: The answer is enlightening! and it uses one of my favorite thing of mathematics. Good luck!

OpenStudy (anonymous):

How is it 2?

OpenStudy (experimentx):

misread question.

OpenStudy (experimentx):

doesn't look like a simple counting problem at all.

OpenStudy (anonymous):

25 is the maximum number of green steps I think. 25c1 + ... + 24c25? Maybe

OpenStudy (experimentx):

I think the same ... but not the latter part.

OpenStudy (anonymous):

well, if you want green steps to never succeed each other, then you can only have one green step or no green steps. which leaves 51 possibilities

OpenStudy (anonymous):

then one of those green steps must come after the other

OpenStudy (anonymous):

No, not necessarily.

OpenStudy (anonymous):

After you color them it won't be.

OpenStudy (anonymous):

I mean red is different from green.

OpenStudy (experimentx):

1+50+(50*48)/2+(50*48*46)/3+ ... should be something very similar like it.

OpenStudy (anonymous):

No sorry.

OpenStudy (anonymous):

Take time guys, if you haven't seen it before or you are not familiar to what I am thinking it will take some time.

OpenStudy (anonymous):

How about, 50c1 + (50c2- 49) + (50c3 -48*49) +...(50c25 - 25*26*27*...*49)?lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Nooooo, Sorry.

OpenStudy (shayaan_mustafa):

ok I think probability should be like that if there are x steps then probability for paint steps with red color will be n(x)=n(x-1)+n(x-2) I think. actually i am so weak in probability.

OpenStudy (shayaan_mustafa):

does my formula is correct either? if so then let me progress my calculation. but first let me know about this brain teasing formula guys..

OpenStudy (shayaan_mustafa):

well guys. according to given conditions i think we must use Fibonaci series.

OpenStudy (experimentx):

answer version 1.1: 1 + 50 +((48*47)/2 + 2*48) + (2*48*46+48*47*45/3) + ...

OpenStudy (experimentx):

@FoolForMath what do you say ... am i pretty close??

OpenStudy (accessdenied):

I think Fibonacci series sounds like a good call when I count for 2-5 stairs, I get 3 for 2 stairs, 5 for 3 stairs, 8 for 4 stairs, 13 for 5 stairs, and 21 for 6 stairs

OpenStudy (shayaan_mustafa):

yes. @accessdenied

OpenStudy (karatechopper):

what do u mean green doesnt succeed each other?

OpenStudy (anonymous):

It's Fibonacci series.

OpenStudy (shayaan_mustafa):

for 20 steps 17711 ways are allowed.

OpenStudy (shayaan_mustafa):

similarly calculate for 50 steps.

OpenStudy (anonymous):

@Shayaan_Mustafa you google?!

OpenStudy (anonymous):

well I ma sure you did.

OpenStudy (karatechopper):

wait u said counting?!?!

OpenStudy (karatechopper):

OMG!

OpenStudy (anonymous):

Surprisingly the original problem ask for 20 steps.

OpenStudy (shayaan_mustafa):

google is not opening due to proxy server.

OpenStudy (shayaan_mustafa):

i am using proxy to serve openstudy. coz our area internet is blocked. and this problem i had solved in college 2 years ago. that's why i recall it. and not sure but it was for 20 steps.

OpenStudy (anonymous):

I agree, that the problem is hard if you haven't solved it before.

OpenStudy (shayaan_mustafa):

i am using proxy to serve openstudy. coz our area internet is blocked. and this problem i had solved in college 2 years ago. that's why i recall it. and not sure but it was for 20 steps.

OpenStudy (shayaan_mustafa):

google is not opening due to proxy server.

OpenStudy (anonymous):

Anyways, congrats on solving foo's problem of the day.

OpenStudy (anonymous):

Why can't you use proxy for google?

OpenStudy (anonymous):

Btw which region is banned?

OpenStudy (shayaan_mustafa):

I don't know from where you got these problem every day and they are really brain teasing.

OpenStudy (shayaan_mustafa):

no it is not banned. due to some server problem some house got stuck.

OpenStudy (anonymous):

I have multiple resource :P

OpenStudy (shayaan_mustafa):

lol... not you own brain??? :p

OpenStudy (anonymous):

I am good with remixing problems and making it interesting, and no I don't author all the problems I post here.

OpenStudy (anonymous):

How did you know fibonacci works here?

OpenStudy (anonymous):

Lucas Fibonacci came to home and whispered into my ears. :P

OpenStudy (anonymous):

I didn't knew his first name was Lucas, thanks.

OpenStudy (shayaan_mustafa):

@Ishaan94 there are talking about steps so I make a guess. and accessdenied give it a positive direction.

OpenStudy (anonymous):

Lol, okay I computed for the small n few terms and I can see the patern

OpenStudy (experimentx):

so i can google at last.

OpenStudy (accessdenied):

i broke it into a smaller problem and counted the number of ways you could arrange the painted green numbers for smaller i didnt notice that it was fibonacci until i saw it mentioned. :P

OpenStudy (anonymous):

You can't guess fibonacci series just like that. It's completely unrelated to this topic or maybe just for me. :/

OpenStudy (anonymous):

Fibonacci, Catalan, Prime numbers, Pi are never unrelated to anything ... :P

OpenStudy (shayaan_mustafa):

no... read the condition which FFM has given. then you can guess it. I am electronics engineer. so it is easy to think about brain teasing questions. because i have those circuits in electronics and solved which have take two weeks in their constructions buddy..

OpenStudy (anonymous):

It takes two weeks just to build one circuit? :o

OpenStudy (anonymous):

It takes more than two for some.

OpenStudy (shayaan_mustafa):

yes. coz it was for function generator. and that was a team work of mine and my friend.

OpenStudy (anonymous):

Oh

OpenStudy (shayaan_mustafa):

FFM is right. coz calculation or circuit analysis is not so easy by hand.

OpenStudy (anonymous):

you can always wear gloves though :P

OpenStudy (shayaan_mustafa):

hehehehhe.... noooooOOO lol

OpenStudy (kinggeorge):

The trivial upper bound would be \(2^{50}\) since that's the total number of ways possible. This upper bound can be easily reduced to \[\large 2^{50} - {49\cdot 50 \over 2}\]if the green steps only appear in a single group (when a group has more than one green step).

OpenStudy (anonymous):

OOPS!

OpenStudy (kinggeorge):

But I don't feel like this is the right way.

OpenStudy (anonymous):

Trust your instincts ;)

OpenStudy (anonymous):

Wild guess: \[\sum_{k=0}^{50}(-1)^k\left(\begin{matrix}50 \\ k\end{matrix}\right)(50-k+1)2^{50-k}\] Don't ask me to explain this, it's probably too complicated to be correct anyway?

OpenStudy (anonymous):

I can almost smell Stirling number ?

OpenStudy (anonymous):

I don't know, I've never heard of Stirling numbers before?

OpenStudy (anonymous):

you may like to read the discussions.

OpenStudy (anonymous):

for n = 1 we have f(n) = 2 G or R by adding on another "stair" first consider it red giving us f(n) then consider it green where we have f(n) - (all those beginning with G) we have f(2) = f(1) + f(1) - 1 =3 RR , GR , RG let (all those beginning with G) = g(n) now g(n+1) = f(n) - g(n) (we did this when we considered it beginning with G) f(n) = f(n-1) +f(n-1) -g(n-1) =f(n-1) + f(n-1) -( f(n-2) -g(n-2))) maybe breaking it down like this? im not sure

OpenStudy (anonymous):

i broke it down like i said

OpenStudy (anonymous):

unless i made some silly mistake

OpenStudy (experimentx):

you are genius!!!

OpenStudy (anonymous):

thanks for your compliment :)

OpenStudy (anonymous):

am i right?

OpenStudy (experimentx):

i haven't tried after my method did not work ...

OpenStudy (anonymous):

i think i made an error

OpenStudy (experimentx):

i think i should give a try too.

OpenStudy (anonymous):

got it i messed up on a sign, and also on my A_r and B_r it cancels at the end its the 53rd fibonacci number

OpenStudy (anonymous):

for n = 1 we have f(n) = 2 G or R by adding on another "stair" first consider it red giving us f(n) then consider it green where we have f(n) - (all those beginning with G) we have f(2) = f(1) + f(1) - 1 =3 RR , GR , RG let (all those beginning with G) = g(n) now g(n+1) = f(n) - g(n) (we did this when we considered it beginning with G) f(n) = f(n-1) +f(n-1) -g(n-1) =f(n-1) + f(n-1) -( f(n-2) -g(n-2)) break it down using recursive: f(n)= 2f(n-1) - g(n-1) g(n) = f(n-1) - g(n-1) so we have f(n) = (A_r)f(n-r) + (B_r)(n-r) we get: \[A_{r+1} = 2A_{r} + B_{r}\] \[B_{r+1} = -(A_{r} + B_{r})\] \[A_1 = 2 \] \[B_1 = -1\] a few calculations reveal the Fibonacci pattern here i then showed inductively that A_r is the (r+ 3)th Fibonacci number and B_r is negative the (r+1)th fib number then let n = 50 , let r = n-1, and we already know f(1) and g(1) some cancellations leave us with 53rd fib number